In Minnesota, K-12 substitute teachers who work whenever asked earn an average yearly

                   salary of $6,000 for teaching. If the distribution of salaries is approximately normal with

                   a standard deviation of $1,000, what is the probability that a randomly selected K-12

                   substitute teachers makes the following:

  1. More than $9,000 a year.              b) Between $3,000 and $5,000 a year.
Answer #1

Ans:

Normal distribution:

mean=6000

standard deviation=1000

a)

z=(9000-6000)/1000

z=3

P(z>3)=0.00135

b)

z(3000)=(3000-6000)/1000=-3

z(5000)=(5000-6000)/1000=-1

P(-1<z<-3)=P(z<-1)-P(z<-3)

=0.15866-0.00135=0.1573